OpenStudy (anonymous):

how do you write an equation of the line with the slope of -2 and containing the point (2,-4)?

OpenStudy (tkhunny):

A line in slope-intercept form looks like this... y = mx + b We are given the slope. -2 Now, it looks like this... y = -2x + b We are given a point. (2,-4). What do you suppose we could do with that information?

OpenStudy (anonymous):

well i know that is going to be the y intercept and it the point was (0,-4) the y intercept would be -4 but since there is now a 2 instead of a 0 i have no idea what to do.

OpenStudy (tkhunny):

Not bad. Just substitute, anyway. y = -2x + b (2,-4) suggests x = 2 when y = -4, so -4 = -2(2) + b And we should be able to solve this for 'b'. No?

OpenStudy (anonymous):

so b would equal 0?

OpenStudy (tkhunny):

Looks like it. Good work.

OpenStudy (anonymous):

hey thanks i appreciate it!